spicyboi420:

A salesperson earns $200 a week plus a 4% commission on her sales. Which equation models the relationship between her sales in a given week, s, and her weekly earnings, e?

supie:

I'm guessing \(\rm s=100\%\\c=4\%\) \(c=\frac{4(s)}{100}\) \(c=\frac{4s}{100}\) Weekly earnings, e. \(\it\ e=200+\frac{4s}{100}\)